Maria's age is M. Pat's age is M+3 five years ago. The product of their age was 70. How old is Maria now?
Answer by stanbon(75887)

Maria's age is M. Pat's age is M+3 five years ago. The product of their age was 70. How old is Maria now?
-----------------
Equation:
M(M+3) = 70
M^2+3M-70 = 0
-----
Factor:
(M+10)(M-7) = 0
Positive solution:
M = 7 yrs (Maria's age now)
M+3 = 10 (Pat's age now)
============================
Cheers,
Stan H.
=============